# A Survey on Large Language Models for Code Generation

### v1

@@ 4.10.1. Metrics
where $n$ is the total number of sampled candidate code solutions, $k$ is the number of randomly selected code solutions from these candidates for each programming problem, with $n\geq k$, and $c$ is the count of correct samples within the $k$ selected. Tables 6 and 7 illustrate the performance of contemporary large language models (LLMs) for code generation, measured by the pass@k metric across different values of $k\in\{1,10,100\}$ on the HumanEval and MBPP benchmarks, respectively.
@@ 4.10.2. Human Evaluation
Table 7. The performance comparison of LLMs for code generation on the MBPP (Austin et al., 2021) benchmark, measured by Pass@{1, 10, 100}. For models with various sizes, we report only the largest size version of each model.
Model Size pass@k
$k=1$ $k=10$ $k=100$
GPT-3.5-Turbo (OpenAI, 2022) - 52.2 - -
Claude-3-Opus (Anthropic, 2024) - 89.4 - -
Claude-3-Haiku (Anthropic, 2024) - 80.2 - -
Claude-3-Sonnet (Anthropic, 2024) - 83.6 - -
StarCoder2-Instruct (Yuxiang Wei, 2024) 15.5B 78 - -
CodeGemma (CodeGemma Team et al., 2024) 7B 65.1 - -
StarCoder 2 (Lozhkov et al., 2024) 15B 50.6 - -
phi-2 (Mojan Javaheripi, 2023) 2.7B 64 - -
WaveCoder (Yu et al., 2023) 6.7B 74.9 - -
CodeFuse (Liu et al., 2023a) 34B 61.0 - -
CodeQwen (Team, 2024) 14B 51.4 - -
DeepSeek Coder (Guo et al., 2024) 33B 66.0 - -
Phi-1.5 (Li et al., 2023b) 1.3B 43.5 - -
WizardCoder (Luo et al., 2023) 16B 51.8 - -
StarCoder (Li et al., 2023a) 5.5B 52.7 - -
SantaCoder (Allal et al., 2023) 1.1B 3.65 21.33 41.92
PyCodeGPT (Zan et al., 2022) 110M 9.39 28.37 48.71
PolyCoder (Xu et al., 2022) 2.7B 4.39 17.99 38.17
phi-1 (Gunasekar et al., 2023) 1.3B 55.5 - -
PaLM-Coder (Chowdhery et al., 2023) 540B 47 - -
PaLM (Chowdhery et al., 2023) 540B 36.8 - -
LLaMA (Touvron et al., 2023a) 65B 37.7 - -
LLaMA 2 (Touvron et al., 2023b) 70B 45.4 66.2 83.1
CodeT5+ (Wang et al., 2021) 16B 56.6 - -
InCoder (Fried et al., 2022) 6.7B 21.3 46.5 66.2
GPT-Neo (Black et al., 2021) 2.7B 5.89 23.09 44.26
GPT-J (Wang and Komatsuzaki, 2021) 6B 11.30 35.62 53.63
CodeT5 (Wang et al., 2021) 770M 15.78 38.63 50.35
CodeParrot (Tunstall et al., 2022) 1.5B 1.29 8.66 27.17
Code Llama (Roziere et al., 2023) 34B 55 76.2 86.6
CodeGen-NL (Nijkamp et al., 2022) 16.1B 10.92 38.43 62.76
CodeGen-Multi (Nijkamp et al., 2022) 16.1B 20.94 51.61 70.02
CodeGen-Mono (Nijkamp et al., 2022) 16.1B 35.28 67.32 80.09
CodeGeeX (Zheng et al., 2023c) 13B 24.4 48 -
BLOOM (Le Scao et al., 2023) 1.7B 3.16 14.23 31.38
PanGu-Coder (Christopoulou et al., 2022) 2.6B 23.0 43.60 59.64
CodeGeeX2 (Zheng et al., 2023c) 6B 24.37 47.95 -
Moreover, in an effort to better align large language models (LLMs) with human preferences and intentions, InstructGPT (Ouyang et al., 2022) employs human-written prompts and demonstrations, and model output ranking in the fine-tuning of LLMs using reinforcement learning from human feedback (RLHF). Although similar alignment learning techniques have been applied to code generation, the feedback in this domain typically comes from a compiler or interpreter, which offers execution feedback, rather than from human evaluators. Notable examples include CodeRL (Le et al., 2022), PPOCoder (Shojaee et al., 2023), RLTF (Liu et al., 2023d), and PanGu-Coder2 (Shen et al., 2023). Further information on this topic is available in Section 4.5.
@@ 4.10.3. LLM-as-a-Judge
The powerful instruction-following capabilities of large language models (LLMs) have stimulated researchers to innovatively investigate the potential of LLM-based evaluations. The LLM-as-a-Judge (Zheng et al., 2024a) refers to the application of advanced proprietary LLMs (e.g., GPT4, Gemini, and Claud 3) as proxies for human evaluators. This involves designing prompts with specific requirements to guide LLMs in conducting evaluations, as demonstrated by AlpacaEval (Li et al., 2023c) and MT-bench (Zheng et al., 2024a). This method reduces reliance on human participation, thereby facilitating more efficient and scalable evaluations. Moreover, LLMs can offer insightful explanations for the assigned rating scores, thereby augmenting the interpretability of evaluations (Zhao et al., 2023b).
@@
Nevertheless, the use of LLM-based evaluation for code generation remains relatively underexplored compared with general-purpose LLM. A recent work (Zhuo, 2024) introduces the ICE-Score evaluation metric, which instructs LLM for code assessments. This approach attains superior correlations with functional correctness and human preferences, thereby eliminating the requirement for test oracles or references. As the capabilities of LLM continue to improve, we anticipate seeing more research in this direction.
@@
Despite their scalability and explainability, the effectiveness of LLM-based evaluation is constrained by the inherent limitations of the chosen LLM. Several studies have shown that most LLMs, including GPT-4, suffer from several issues, including position, verbosity, and self-enhancement biases, as well as restricted reasoning ability (Zheng et al., 2024a). Specifically, position bias refers to the tendency of large language models (LLMs) to disproportionately favor responses that are presented in certain positions, which can skew the perceived quality of answers based on their order of presentation. Meanwhile, verbosity bias describes the inclination of LLMs to prefer lengthier responses, even when these are not necessarily of higher quality compared to more concise ones. Self-enhancement bias, on the other hand, is observed when LLMs consistently overvalue the quality of the text they generate (Zheng et al., 2024a; Zhao et al., 2023b). Moreover, due to their inherent limitations in tackling complex reasoning challenges, LLMs may not be entirely reliable as evaluators for tasks that require intensive reasoning, such as those involving mathematical problem-solving. However, these shortcomings can be partially addressed through the application of deliberate prompt engineering and fine-tuning techniques, as suggested by (Zheng et al., 2024a).

### v2

@@ 5.10.1. Metrics
where $n$ is the total number of sampled candidate code solutions, $k$ is the number of randomly selected code solutions from these candidates for each programming problem, with $n\geq k$, and $c$ is the count of correct samples within the $k$ selected.
@@ 5.10.2. Human Evaluation
Moreover, in an effort to better align LLMs with human preferences and intentions, InstructGPT (Ouyang et al., 2022) employs human-written prompts and demonstrations, and model output ranking in the fine-tuning of LLMs using reinforcement learning from human feedback (RLHF). Although similar alignment learning techniques have been applied to code generation, the feedback in this domain typically comes from a compiler or interpreter, which offers execution feedback, rather than from human evaluators. Notable examples include CodeRL (Le et al., 2022), PPOCoder (Shojaee et al., 2023), RLTF (Liu et al., 2023e), and PanGu-Coder2 (Shen et al., 2023). Further information on this topic is available in Section 5.5.
@@ 5.10.3. LLM-as-a-Judge
Figure 16. The pipeline of (Code) LLM-as-a-judge for evaluating generated code by Code LLMs. There are primarily two types of approaches: pairwise comparison and single answer grading.
@@
The powerful instruction-following capabilities of LLMs have stimulated researchers to innovatively investigate the potential of LLM-based evaluations. The LLM-as-a-Judge (Zheng et al., 2024a) refers to the application of advanced proprietary LLMs (e.g., GPT4, Gemini, and Claud 3) as proxies for human evaluators. This involves designing prompts with specific requirements to guide LLMs in conducting evaluations, as demonstrated by AlpacaEval (Li et al., 2023c) and MT-bench (Zheng et al., 2024a). This method reduces reliance on human participation, thereby facilitating more efficient and scalable evaluations. Moreover, LLMs can offer insightful explanations for the assigned rating scores, thereby augmenting the interpretability of evaluations (Zhao et al., 2023b).
@@
Nevertheless, the use of LLM-based evaluation for code generation remains relatively underexplored compared with general-purpose LLM. The pipeline of (Code) LLM-as-a-judge for evaluating generated code by Code LLMs is depicted in Figure 16. A recent work (Zhuo, 2024) introduces the ICE-Score evaluation metric, which instructs LLM for code assessments. This approach attains superior correlations with functional correctness and human preferences, thereby eliminating the requirement for test oracles or references. As the capabilities of LLM continue to improve, we anticipate seeing more research in this direction.
@@
Despite their scalability and explainability, the effectiveness of LLM-based evaluation is constrained by the inherent limitations of the chosen LLM. Several studies have shown that most LLMs, including GPT-4, suffer from several issues, including position, verbosity, and self-enhancement biases, as well as restricted reasoning ability (Zheng et al., 2024a). Specifically, position bias refers to the tendency of LLMs to disproportionately favor responses that are presented in certain positions, which can skew the perceived quality of answers based on their order of presentation. Meanwhile, verbosity bias describes the inclination of LLMs to prefer lengthier responses, even when these are not necessarily of higher quality compared to more concise ones. Self-enhancement bias, on the other hand, is observed when LLMs consistently overvalue the quality of the text they generate (Zheng et al., 2024a; Zhao et al., 2023b). Moreover, due to their inherent limitations in tackling complex reasoning challenges, LLMs may not be entirely reliable as evaluators for tasks that require intensive reasoning, such as those involving mathematical problem-solving. However, these shortcomings can be partially addressed through the application of deliberate prompt engineering and fine-tuning techniques, as suggested by (Zheng et al., 2024a).
5.10.4. Empirical Comparison
Figure 17. The performance comparison of LLMs for code generation on the MBPP (Austin et al., 2021) benchmark, measured by Pass@1. For models with various sizes, we report only the largest size version of each model with a magnitude of B parameters.
Figure 18. The performance comparison of LLMs for code generation on the BigCodeBench (Zhuo et al., 2024a) benchmark, measured by Pass@1. For models with various sizes, we report only the largest size version of each model with a magnitude of B parameters.
In this section, we present a performance comparison of LLMs for code generation using the well-regarded HumanEval, MBPP, and the more practical and challenging BigCodeBench benchmarks. This empirical comparison aims to highlight the progressive enhancements in LLM capabilities for code generation. These benchmarks assess an LLM’s ability to generate source code across various levels of difficulty and types of programming tasks. Specifically, HumanEval focuses on complex code generation, MBPP targets basic programming tasks, and BigCodeBench emphasizes practical and challenging programming tasks.
Due to the limitations in computational resources we faced, we have cited experimental results from original papers or widely recognized open-source leaderboards within the research community, such as the HumanEval Leaderboard  [^16] https://paperswithcode.com/sota/code-generation-on-humaneval, EvalPlus Leaderboard  [^17] https://evalplus.github.io/leaderboard.html, Big Code Models Leaderboard  [^18] https://huggingface.co/spaces/bigcode/bigcode-models-leaderboard, and BigCodeBench Leaderboard  [^19] https://bigcode-bench.github.io/. We report performance on HumanEval using the pass@1 metric, as shown in Table 9, while MBPP and BigCodeBench results are presented with pass@1 in Figures 17 and 18, respectively.
We offer the following insights:
  • The performance gap between open-source and closed-source models across the three benchmarks is gradually narrowing. For instance, on the HumanEval benchmark, DeepSeek-Coder-V2-Instruct with 21B activation parameters and Qwen2.5-Coder-Instruct 7B achieve 90% and 88.4% pass@1, respectively. These results are comparable to the much larger closed-source LLMs, such as Claude-3.5-Sonnet, which achieves 92.0% pass@1. On the MBPP benchmark, Qwen2.5-Coder-Instruct 7B with 83.5% pass@1 significantly outperforms GPT-3.5-Turbo with 52.2% pass@1 and closely rivals the closed-source Claude-3-Opus with 86.4% pass@1. On the BigCodeBench, DeepSeek-Coder-V2-Instruct achieves 59.7%, surpassing all compared closed-source and open-source LLMs except for slightly falling behind GPT-4o-0513, which achieves 61.1%.
  • Generally, as the number of model parameters increases, the performance of code LLMs improves. However, Qwen2.5-Coder-Instruct 7B achieves 88.4% pass@1, outperforming larger models like StarCoder2-Instruct 15.5B with 72.6% pass@1, DeepSeek-Coder-Instruct 33B with 79.3% pass@1, and Code Llama-Instruct 70B with 67.8% pass@1 on the HumanEval benchmark. Similar trends are observed across the other two benchmarks, suggesting that code LLMs with 7B parameters may be sufficiently capable for code generation task.
  • Instruction-tuned models consistently outperform their base (pretrained) counterparts across the HumanEval and MBPP benchmarks. For instance, Qwen2.5-Coder-Instruct surpasses Qwen2.5-Coder by an average of 26.04%, StarCoder2-Instruct improves upon StarCoder 2 by an average of 35.20%, and CodeGemma-Instruct enhances CodeGemma by an average of 11.26%. Additionally, DeepSeek-Coder-Instruct outperforms DeepSeek-Coder by an average of 23.71%, while Code Llama-Instruct shows a 13.80% improvement over Code Llama. Detailed results can be found in Table 10. These findings underscore the effectiveness of instruction tuning, although the quality of the instruction tuning dataset plays a critical role in determining model performance (Luo et al., 2023; Zhou et al., 2024).
    Table 10. The performance improvement of instruction-tuned models over their pretrained counterparts on the HumanEval, MBPP, and BigCodeBench benchmarks. The last two rows demonstrate the average improvement on the first two benchmarks and the three benchmarks, respectively.
    Qwen2.5-Coder- Instruct 7B StarCoder2- Instruct 15.5B CodeGemma- Instruct 7B DeepSeek-Coder- Instruct 33B Code Llama- Instruct 70B
    HumanEval 43.51% 56.80% 26.07% 41.35% 27.92%
    MBPP 8.58% 13.60% -3.56% 6.06% -0.32%
    BigCodeBench - 17.45% 2.61% 9.66% 12.73%
    # Avg. Imp. H. M. 26.04% 35.20% 11.26% 23.71% 13.80%
    # Avg. Imp. H. M. B. - 29.28% 8.37% 19.02% 13.44%
  • Performance on the HumanEval benchmark is nearly saturated. However, MBPP, which involves basic programming tasks, and BigCodeBench, which involves more practical and challenging programming tasks, demand more capable code LLMs. Additionally, while these benchmarks primarily evaluate the functional correctness of code, they do not provide a comprehensive assessment across other critical dimensions. Developing a more holistic evaluation framework that integrates various aspects remains an open area for future research and development in LLMs for code generation evaluation.
Discussion: We discuss certain code LLMs in Table 9 for clarity: (1) General LLMs accessed via API are not specifically trained on large code corpora but achieve state-of-the-art performance in code generation, such as Claude-3.5-Sonnet with 92.0% pass@1 on HumanEval benchmark. (2) AlphaCode targets code generation for more complex and unseen problems that require a deep understanding of algorithms and intricate natural language, such as those encountered in competitive programming. The authors of AlphaCode found that large-scale model sampling to navigate the search space, such as 1M samples per problem for CodeContests, followed by filtering based on program behavior to produce a smaller set of submissions, is crucial for achieving good and reliable performance on problems that necessitate advanced reasoning. (3) Phi-1 1.3B is a specialized LLM for code, trained on “textbook quality” data from the web (6B tokens) and synthetically generated textbooks and exercises using GPT-3.5 (1B tokens). (4) Code Llama 70B is initialized with Llama 2 model weights and continually pre-trained on 1T tokens from a code-heavy dataset and long-context fine-tuned with approximately 20B tokens. However, Code Llama-Instruct 70B is fine-tuned from Code Llama-Python 70B without long-context fine-tuning, using an additional 260M tokens to better follow human instructions. Surprisingly, these models underperform compared to smaller parameter Code LLMs like Qwen2.5-Coder-Instruct 7B, DeepSeek-Coder-V2-Instruct 21B, and Codestral 22B across all three benchmarks. The underlying reasons for this discrepancy remain unclear and warrant further exploration. (5) Unlike other open-source Code LLMs, DeepSeek-Coder-V2-Instruct is further pre-trained on DeepSeek-V2 (Liu et al., 2024a), which employs a Mixture-of-Experts (MoE) architecture with only 21B activation parameters out of 236B parameters, using an additional 6 trillion tokens composed of 60% source code, 10% math corpus, and 30% natural language corpus. For a comprehensive understanding of MoE in LLMs, please refer to (Cai et al., 2024b).
5.11. Code LLMs Alignment
The pre-training of LLMs for next-token prediction, aimed at maximizing conditional generation likelihood across vast textual corpora, equips these models with extensive world knowledge and emergent capabilities (Brown, 2020). This training approach enables the generation of coherent and fluent text in response to diverse instructions. Nonetheless, LLMs can sometimes misinterpret human instructions, produce biased content, or generate factually incorrect information (commonly referred to as hallucinations), which may limit their practical utility (Wang et al., 2023c; Zhao et al., 2023b; Ji et al., 2023b).
Aligning LLMs with human intentions and values, known as LLM alignment, has consequently become a critical research focus (Ji et al., 2023b; Wang et al., 2023c). Key objectives frequently discussed in the context of LLM alignment include robustness, interpretability, controllability, ethicality, trustworthiness, security, privacy, fairness, and safety. In recent years, significant efforts have been made by researchers to achieve this alignment, employing techniques such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022).
However, the alignment of Code LLMs has not been extensively explored. Compared to text generation, aligning code generation with human intentions and values is even more crucial. For instance, users without programming expertise might prompt Code LLM to generate source code and subsequently execute it on their computers, potentially causing catastrophic damage. Some potential risks include:
  • Malware Infection: The code could contain viruses, worms, or trojans that compromise our system’s security.
  • Data Loss: It might delete or corrupt important files and data.
  • Unauthorized Access: It can create backdoors, allowing attackers to access our system remotely.
  • Performance Issues: The code might consume excessive resources, slowing down our system.
  • Privacy Breaches: Sensitive information, such as passwords or personal data, might be stolen.
  • System Damage: It may alter system settings or damage hardware components.
  • Network Spread: It could propagate across networks, affecting other devices.
  • Financial Loss: If the code is ransomware, it might encrypt data and demand payment for decryption.
  • Legal Consequences: Running certain types of malicious code can lead to legal repercussions.
As illustrated, aligning Code LLMs to produce source code consistent with human preferences and values is of paramount importance in software development. A recent study (Yang et al., 2024a) provides the first systematic literature review identifying seven critical non-functional properties of LLMs for code, beyond accuracy, including robustness, security, privacy, explainability, efficiency, and usability. This study is highly pertinent to the alignment of Code LLMs. We recommend readers refer to this survey for more detailed insights.
In this survey, we identify five core principles that serve as the key objectives for aligning Code LLMs: Green, Responsibility, Efficiency, Safety, and Trustworthiness (collectively referred to as GREST). These principles are examined from a broader perspective. Each category encompasses various concepts and properties, which are summarized in Table 11.
Table 11. Five core principles serve as the key objectives for Code LLMs alignment: Green, Responsibility, Efficiency, Safety, and Trustworthiness (collectively referred to as GREST).
Principles Involved Concepts and Properties
Green Energy Efficiency: Minimizing computational energy use and reduce environmental impact and financial costs. Sustainable Materials: Leveraging eco-friendly infrastructure and servers for code generation, lowering long-term expenses. Carbon Footprint: Reducing emissions associated with model training and inference to enhance efficiency and save costs. Resource Optimization: Efficiently utilizing computational resources to minimize waste and reduce expenses in code generation. Recycling Management: Responsibly dispose of hardware used in model development to reduce waste management costs. Renewable Energy: Utilizing renewable energy sources for powering training and inference processes to decrease energy costs. Lifecycle Assessment: Evaluating the environmental and financial impacts of models from creation to deployment and disposal.
Responsibility Ethical Considerations: Adhering to ethical guidelines to ensure responsible use and deployment of generated code. Accountability: Establishing clear lines of responsibility for code generation outcomes and potential impacts. User Education: Providing resources and guidance to help users understand and responsibly use generated code. Impact Assessment: Evaluating the social and technical implications of code generation to minimize negative effects. Regulatory Compliance: Ensuring that generated code adheres to relevant laws (e.g., copyright) and industry regulations.
Efficiency Model Optimization: Streamlining models to reduce computational load and improve speed. Prompt Engineering: Designing effective prompts to generate accurate code efficiently. Resource Management: Allocating computational resources wisely to balance speed and cost. Inference Optimization: Enhancing the inference process to quickly generate code with minimal latency. Parallel Processing: Utilizing parallelism to speed up code generation tasks. Caching Mechanisms: Implementing caching to reuse previous results and reduce redundant computations. Evaluation Metrics: Using precise metrics to assess and improve the efficiency of code outputs.
Safety Input Validation: Ensuring inputs (prompts) are safe and sanitized to prevent malicious exploitation. Security Audits: Regularly reviewing generated code for vulnerabilities and potential exploits. Monitoring and Logging: Keeping track of generation outputs to quickly identify and address safety issues. User Access Control: Limiting access to generation capabilities to trusted users to minimize risk. Continuous Updates: Regularly updating models with the latest safety protocols and security patches. Ethical Guidelines: Implementing ethical standards to guide safe and responsible code generation.
Trustworthiness Reliability: Ensuring that generated code consistently meets functional requirements and performs as expected. Transparency: Providing clear explanations of how code is generated to build user confidence. Verification and Testing: Using rigorous testing frameworks to ensure the generated code accuracy and reliability. Bias Mitigation: Actively working to identify and reduce biases in code generation to ensure fairness and impartiality. User Feedback Integration: Continuously incorporating user feedback to refine and improve code generation processes. Documentation: Providing comprehensive documentation for generated code to enhance understanding and trust.
In the following, we define each principle and briefly introduce a few notable works to enhance understanding.
Green: The Green principle underscores the importance of environmental sustainability in the development and deployment of LLMs for code generation. This involves optimizing energy consumption and reducing both the carbon footprint and financial costs associated with training and inference processes. Currently, training, inference, and deployment of Code LLMs are notably resource-intensive. For example, training GPT-3, with its 175 billion parameters, required the equivalent of 355 years of single-processor computing time and consumed 284,000 kWh of energy, resulting in an estimated 552.1 tons of CO2 emissions (Samsi et al., 2023). Furthermore, a ChatGPT-like application, with an estimated usage of 11 million requests per hour, can produce emissions of 12.8k metric tons of CO2 per year, which is 25 times the carbon emissions associated with training GPT-3 (Chien et al., 2023). To mitigate these costs, several techniques are often employed, such as the development of specialized hardware (e.g., Tensor Processing Units (TPUs) and Neural Processing Units (NPUs)), model compression methods (e.g., quantization and knowledge distillation), parameter-efficient fine-tuning (PEFT), and the use of renewable energy sources. For instance, Shi et al. (Shi et al., 2024) applied knowledge distillation to reduce the size of CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), resulting in optimized models of just 3MB. These models are 160 times smaller than the original large models and significantly reduce energy consumption by up to 184 times and carbon footprint by up to 157 times. Similarly, Wei et al. (Wei et al., 2023a) utilized quantization techniques for Code LLMs such as CodeGen (Nijkamp et al., 2022) and Incoder (Fried et al., 2022) by employing lower-bit integers (e.g., int8). This approach reduced storage requirements by 67.3% to 70.8%, carbon footprint by 28.8% to 55.0%, and pricing costs by 28.9% to 55.0%.
Responsibility: The Responsibility principle in the context of Code LLMs underscores the importance of ethical considerations, fairness, and accountability throughout their lifecycle. This involves addressing biases in training data, ensuring fairness and transparency in model decision-making, maintaining accountability for outputs, adhering to applicable laws (e.g., copyright), implementing safeguards against misuse, and providing clear communication about the model’s capabilities and limitations. Specifically,
  • Bias Mitigation. Biases in code generation can lead to flawed software and reinforce stereotypes, potentially causing significant societal impacts. For example, an Code LLM that inherits biases from its training data may produce source code/software that inadvertently discriminates against certain user groups. This can result in applications that fail to meet the diverse needs of users, promoting exclusionary practices and reinforcing existing stereotypes (Mouselinos et al., 2022; Liu et al., 2023a).
  • Fairness and Transparency. A lack of fairness and transparency in Code LLM decision-making can result in biased or suboptimal code solutions. If the model’s decision-making process is opaque, developers might unknowingly introduce code that favors specific frameworks or libraries, thereby limiting innovation and diversity in software development. This opacity can create unfair advantages and hinder collaborative efforts within tech communities (Bogina et al., 2022).
  • Legal Compliance. Compliance with relevant laws, such as licensing and copyright, is crucial when using Code LLMs for code generation to avoid legal complications. If an Code LLM generates code snippets that inadvertently infringe on existing copyrights, it can lead to legal disputes and financial liabilities for developers and organizations (Xu et al., 2024b). Such risks may discourage the use of advanced AI tools, thus stifling innovation and affecting growth and collaboration within the tech community.
  • Accountability. Without accountability for code generated by Code LLMs, addressing bugs or security vulnerabilities becomes challenging. If a model generates faulty code leading to a security breach, the absence of clear accountability can result in significant financial and reputational damage for companies. This uncertainty can delay critical issue resolution and impede trust in AI-assisted development (Liesenfeld et al., 2023).
  • Misuse Prevention. Failing to implement mechanisms to prevent the misuse of Code LLMs can enable the creation of harmful software. For example, models could be exploited to generate malware or unauthorized scripts, posing cybersecurity risks. Without proper safeguards, these models can facilitate malicious activities, threatening both individual and organizational security (Mousavi et al., 2024).
  • Clear Communication. Without clear communication about a model’s capabilities and limitations, developers may misuse the model or overestimate its abilities. Relying on the model to generate complex, mission-critical code without human oversight can lead to significant software failures. Misunderstanding its limitations can result in faulty implementations and lost productivity (Ross et al., 2023).
To adhere to this principle, potential mitigation methods include bias detection and mitigation, quantification and evaluation, and adherence to ethical guidelines. Liu et al. (Liu et al., 2023a) propose a new paradigm for constructing code prompts, successfully uncovering social biases in code generation models, and developing a dataset along with three metrics to evaluate overall social bias. Recently, Xu et al. (Xu et al., 2024b) introduced LiCoEval, an evaluation benchmark for assessing the license compliance capabilities of LLMs. Additionally, incorporating diverse perspectives in development teams and engaging with stakeholders from various communities can further align Code LLM outputs with ethical standards and societal values.
Efficiency: The Efficiency principle emphasizes optimizing the performance and speed of Code LLMs for code generation while minimizing the computational resources required for training and inference. For instance, training the GPT-3 model, which consists of 175 billion parameters, demands substantial resources. It requires approximately 1,024 NVIDIA V100 GPUs, costing around $4.6$ million and taking approximately 34 days to complete the training process. To address these challenges, various techniques are employed, including model compression methods such as pruning, quantization, and knowledge distillation. Additionally, optimized algorithms like AdamW, parallel strategies such as tensor, pipeline, and data parallelism, and parameter-efficient fine-tuning (PEFT) (see Section 5.4.2) are often utilized. For a comprehensive and detailed discussion on methods to enhance the efficiency of Code LLMs for code generation, please refer to Section 4.5.2, “Efficiency Enhancement”, in (Yang et al., 2024a).
Safety: The Safety principle of Code LLMs is of utmost importance due to their potential to introduce vulnerabilities, errors, or privacy breaches into software systems. Ensuring safety involves comprehensive testing and validation processes to detect and mitigate these risks. For instance, attackers might compromise the training process of LLMs by injecting malicious examples into the training data, a method known as data poisoning attacks (Schuster et al., 2021). Even when attackers lack access to the training process, they may employ techniques like the black-box inversion approach introduced by Hajipour et al. (Hajipour et al., 2024). This method uses few-shot prompting to identify prompts that coax black-box code generation models into producing vulnerable code. Furthermore, Yang et al. (Yang et al., 2024b) and Al-Kaswan et al. (Al-Kaswan et al., 2024) reveals that Code LLMs, such as CodeParrot (Face, 2023), can memorize training data, potentially outputting personally identifiable information like emails, names, and IP addresses, thereby posing significant privacy risks. Additionally, Yuan et al. (Yuan et al., 2023a) demonstrate that engaging with ChatGPT and GPT-4 in non-natural languages can circumvent safety alignment measures, leading to unsafe outcomes, such as “The steps involved in stealing money from a bank.”. To bolster the safety of LLMs in code generation, it is crucial to detect and eliminate privacy-related information from training datasets. For example, approaches outlined in (Fried et al., 2022) and (Allal et al., 2023) utilize carefully crafted regular expressions to identify and remove private information from training data. To counteract black-box inversion, implementing prompt filtering mechanisms is recommended to identify and block prompts that might result in insecure code generation. Moreover, adversarial training can enhance the model’s resilience to malicious prompts. Employing reinforcement learning methods can further align Code LLMs with human preferences, thereby reducing the likelihood of producing harmful outputs.
Trustworthiness: The Trustworthiness principle focuses on developing Code LLMs that users can depend on for accurate and reliable code generation, which is crucial for their acceptance and widespread adoption. Achieving this requires ensuring model transparency, providing explanations for decisions, and maintaining consistent performance across various scenarios. For instance, Ji et al. (Ji et al., 2023a) propose a causal graph-based representation of prompts and generated code to identify the causal relationships between them. This approach offers insights into the effectiveness of Code LLMs and assists end-users in understanding the generation. Similarly, Palacio et al. (Palacio et al., 2023) introduce ASTxplainer, a tool that extracts and aggregates normalized model logits within Abstract Syntax Tree (AST) structures. This alignment of token predictions with AST nodes provides visualizations that enhance end-user understanding of Code LLM predictions. Therefore, by prioritizing trustworthiness, we can bolster user confidence and facilitate the integration of Code LLMs into diverse coding environments. By adhering to the aforementioned principles as key objectives for aligning Code LLMs, researchers and developers can create LLMs for code generation that are not only capable but also ethical, sustainable, and user-centric.